ralph chase plans to sell a piece of property for $180,000. He wants the money to be paid off in two ways- a short term note at 9% interest and a long term at 7% interest. find the amount of each note if the total annual interest paid is $14,600.
Answer by Maths68(1474)



Total Amount = $180,000
Total Interest = $14,600
Amount to be paid off at 9% = x
Amount to be paid off at 7% = 180,000 - x
Amount to be paid off at 9% * 9% + Amount to be paid off at 7% * 7% = Total Interest
x*9/100 + (180000-x)*7/100=14600
x*0.09+(180000-x)*0.07=14600
0.09x+12600-0.07x=14600
0.02x=14600-12600
0.02x=2000
x=2000/0.02
x=100,000

Amount to be paid off at 9% = x = $100,000
Amount to be paid off at 7% = 180,000 - 100,000 = $80,000

Check
======
100000 * 0.09 + 80000 * 0.07 = 14600
9000+5600=14600
14600=14600